What’s the average velocity of a bicycler who travels 4567 meters north in 15 minutes
Answer by Theo(13342)

here's a definition from wikipedia.

From Wikipedia, the free encyclopedia. Velocity is a measure of how fast something moves in a particular direction. To define it needs both magnitude and direction. If an object moves east at 9 meters per second (9m/s), then its velocity is 9 m/s to the east.

in your problem, the bicycler travels 4567 meters north in 15 minutes.
that would be a velocity of 4567 meters / 15 = 304.466666667 meters per minute.
that would be exactly 304 + 7/15 meters per minute.
15 * (304 + 7/15) = 15 * 304 + 15 * 7/15 = 4560 + 7 = 4567 meters.

note that velocity and speed are not the same.
speed doesn't depend on the direction of travel.
velocity does.

here's a fairly decent reference on the difference between speed and velocity.

https://www.mathsisfun.com/measure/speed-velocity.html